### v1

# The Great Comets of 1843 and 1882 at Their Previous Return to Perihelion in the Twelfth Century: One Spectacular, the Other Dull

Abstract
New insights into the history of C/1843 D1 and C/1882 R1, the two celebrated Kreutz sungrazers, are provided by assessing evidence on their appearance at the previous perihelion return, known as X/1106 C1 and the Chinese comet of 1138 (Ho’s No. 403), respectively. The conditions differed vastly because of disparities in geocentric distance, solar elongation, and phase correction (forward scattering), all linked to the arrival times (early February vs early August). The conclusions include: the daytime sighting of the 1106 comet by Sigebert de Gembloux is consistent with expectation and so are the accounts of an exceptionally long tail observed later in twilight; the comet reached perihelion only hours before its daytime detection; the 1138 comet could have never been sighted in daylight or discovered much earlier than it actually was, in early September, one month after perihelion; at discovery, the tail is predicted to have reached elevations of ${}^{\circ}\!$, while the head was merely $\sim$10 above horizon, when observed from moderate northern latitudes; notion that Kreutz sungrazers at perihelion between mid-May and mid-August could not be seen from the ground except possibly in daylight is misleading; the appearance of the 1138 comet’s nucleus after its tidal fragmentation at perihelion is modeled on the assumption that it consisted of five major fragments (including C/1882 R1 and C/1965 S1); and unpredictable morphological changes with time in the 1882 sungrazer’s split nucleus are discussed.
Subject headings:
comets: X/1106 C1, 1138, X/1702 D1, C/1882 R1, C/1965 S1; methods: data analysis
slugcomment: Version May 20, 2025
1. Introduction
Even though the perception of the Kreutz sungrazer system has over the past decades been changing profoundly, one attribute has remained untouched for nearly sixty years, since Marsden’s (1967) pioneering work: regardless of the number of populations (or subgroups) that the system is believed to consist of, the two preeminent ones, incorporating most observed members, have persisted. Marsden called them Subgroups I and II, and they were the only ones known in the 1960s. Significantly, of the two most memorable sungrazers, both appearing in the 19th century less than 40 years apart, one was a member of Subgroup I (now also called Population I) and the other a member of Subgroup II (or Population II). I do of course refer to the Great March Comet of 1843 (C/1843 D1) and the Great September Comet of 1882 (C/1882 R1).
The remarkable orbital work by Hubbard (1851, 1852) demonstrated that the orbital period of the 1843 sungrazer was confined to a likely range of 500–800 years. And the monumental effort by Kreutz (1888, 1891) made it clear that the orbital periods of the four nuclei, the most prominent products of the 1882 sungrazer’s perihelion fragmentation, ranged between 650 and 950 years. When comet Ikeya-Seki (C/1965 S1), another member of Population II, arrived in the practically identical orbit 83 years following the 1882 sungrazer and Marsden (1967) determined its barycentric orbital period to equal about 850 years, it was likely that (i) the two objects had separated from one another early in the 12th century and (ii) the orbital period of nucleus B (or No. 2 in Kreutz’s notation) of the 1882 sungrazer, of about 770 years, almost coincided with the orbital period of the comet before it split at perihelion. The union of the 1882 sungrazer and Ikeya-Seki in a single body prior to their 12th-century perihelion was proven by Marsden rigorously, via direct orbit integration; nucleus B was obviously the largest mass of the 1882 sungrazer.
Given that sungrazers have been appearing over many centuries, even millennia, the gap of less than 40 years between the perihelion times of the presumably most massive members of the Kreutz system looks suspiciously small to say the least. This near coincidence makes one speculate that — in analogy to the Population II pair — the gap could be an accumulated effect of the rate of their separation in the past. However, compared to the 1882 sungrazer and Ikeya-Seki, two major discrepancies become apparent: (i) the small gap accumulated over the time shows that the orbital periods of the 1843 and 1882 comets were very similar, whereas (ii) their other orbital elements differed much more. These arguments prompt one to conclude that the putative breakup of the progenitor comet into its two massive survivors was characterized by an orbital-momentum exchange with a significant contribution in the direction normal to the orbital plane and an orbital-plane contribution that affected appreciably the perihelion distance but only marginally the orbital period. It turns out that these conditions are satisfied only if the fragmentation event took place at large heliocentric distance.
The bottom line is that the 1843 and 1882 sungrazers must have very similar orbital periods, having passed their previous perihelion in the early 12th century in the same order and at times closer to one another than in the 19th century. Since only one suspect in the early 12th century had until recently been known — the Great Comet of 1106 (X/1106 C1) — we were confronted with the problem of a missing second sungrazer.
2. Search for the Missing Sungrazer:
Past Investigations
An important orbital property common to all Kreutz sungrazers is their fixed direction of the line of apsides, except for a minor difference of about ${}^{\circ}\!$.5 between Populations I and II. For example, the line of apsides of the Great March Comet of 1843 is defined by the ecliptic longitude and latitude of perihelion equaling, respectively, ${}^{\circ}\!$.58 and ${}^{\circ}\!$.29 (equinox J2000), showing that the comet arrived from the direction of the constellation of Canis Major, a point a few degrees north of Sirius.
It is well known that, in general, the southern hemisphere is favored and that the observed performance of a sungrazer of the given intrinsic brightness depends strongly on the angle that the line of apsides makes with the Sun-Earth line at the time the comet is at perihelion, implying significant seasonal variations. Because essentially all preperihelion ejecta are sublimated away at close proximity of perihelion, sungrazers become spectacular objects along the post-perihelion branch of the orbit because of their extremely high rates of dust emission primarily in the first hours after perihelion. An intrinsically bright sungrazer makes a spectacle in the evening sky when its perihelion occurs in January–March, in the morning sky when in September–November. December arrivals reach the southernmost areas of the sky.
Problematic is the appearance of Kreutz sungrazers passing perihelion at “wrong” times of the year. While I do not fully subscribe to it, the verdict found in most cometary publications is closely matched by Marsden’s (1967) narrative that a sungrazer “at perihelion between mid-May and mid-August will undoubtedly be missed, unless it can be seen in daylight. The geometry is such that the comet would approach and leave the sun from behind, and it would never be seen in a dark sky.”
One of the issues that I am having with this widely accepted notion involves the comet of 1041 in Hasegawa & Nakano’s (2001) list of suspected sungrazers. This object, observed in China and Korea from September 1 on, was estimated by the authors to have passed perihelion on August 4 $\pm$ 4, more than a week before the end of the disadvantaged time zone. No daylight sightings were reported, but the comet was observed about four weeks after perihelion, when Hasegawa & Nakano computed its location to have been some 1.7 AU from the Earth, 1 AU from the Sun, and a solar elongation of 35.
When Sekanina & Kracht’s (2022) in-depth investigation of the motion of Ikeya-Seki (C/1965 S1) showed that the comet previously was at perihelion around the year 1140 and the Great Comet of 1106 (X/1106 C1) could not be related to either Ikeya-Seki or the Great September Comet of 1882, perfunctory search in Ho’s (1962) catalogue resulted in the discovery of a Chinese comet of 1138, under No. 403, as the likely candidate. Entirely by coincidence, its timeline strongly resembled the timeline of the comet of 1041: the first observation on September 3 and an estimated perihelion time on August 1.
Detailed examination of the conditions for the likely obsevation site indicated that at midpoint of the astronomical twilight, about 4:30 local time on 1138 September 3, the comet’s head would have been located, as reported, in the east. Its predicted elevation was about 11, its solar elongation 36, and it was nearly 1 mag brighter than a computed limiting magnitude of 3.2 for the naked eye; the brightest part of the dust tail was expected to be approximately 8 long at position angle near 270, pointing up at an angle with the vertical, its end at least 15 above the horizon. To the naked eye it was a rather unimpressive object, even though the adopted intrinsic brightness on its approach to perihelion equaled that of comet X/1106 C1. There was no report of daylight sighting, again paralleling the case of the 1041 comet.
Reports of the detections of these two objects in a dark sky (or nearly dark sky of the astronomical twilight) and no daylight detection reports appear to contradict the above generally accepted viewpoint of the Kreutz sungrazers at perihelion in mid-May through mid-August, and this controversy needs to be examined.
Although the comet of 1041 is of major interest as a possible previous appearance of comet Pereyra (Sekanina & Kracht 2022), the primary subject of this paper is the Chinese comet of 1138 as the previous appearance of the Great September Comet of 1882 and comet Ikeya-Seki. In the following I compare the 1138 comet with the Great Comet of 1106; this is particularly appropriate because both objects are believed to have been of about equal intrinsic brightness, yet the 1106 comet was seen in broad daylight $\sim$1 from the Sun on February 2 and displayed a spectacular tail up to $\sim$100 in the following weeks (e.g., Kronk 1999), while the 1138 comet was a lackluster object, reported for the first time more than 30 days after the estimated perihelion time. Yet, the 1106 comet disappeared some 45–50 days after perihelion, while the 1138 comet about 60 days after perihelion. This effect may have been a by-product of the 1138 comet’s more extensive perihelion fragmentation, which is believed to lead to a slower rate of post-perihelion fading (Sekanina 2022; Section 3.1).
3. Comets of 1106 and 1138 Shortly After Passing
Perihelion, and the Limiting Magnitude
The part of a bright comet that has the greatest chance to be seen with the naked eye in broad daylight is the head because of its high surface brightness. In twilight and at night, the most obvious part is the tail because of its extent and greater angular distance from the Sun in the sky. To compare the positions, relative to the Sun and Earth, of the heads and tails of the two comets, I begin with their adopted orbital elements, presented in Table 1. The set for X/1106 C1 comes from Sekanina & Kracht’s (2022) integration of the motion of the Great March Comet of 1843; for the Chinese comet of 1138 from their integration of the motion of nucleus B of the Great September Comet of 1882. The perihelion time of the 1106 comet, derived by Hasegawa & Nakano (2001), was adjusted to better fit the early observations.
Next, I derive the conditions under which observations of the two comets’ heads were, or would have been, made over the first ten days after perihelion. For either comet, the columns in Table 2 provide at each time the heliocentric and geocentric distances, apparent position relative to the Sun in the sky in the polar coordinates (solar elongation and position angle), and phase angle. [^1] Even though an adopted orbit for the 1106 comet is presented in Table 1 and used in Table 2, the object is still designated with X/, since no orbit derived from the 1106 observations is available.
Table 2 demonstrates that the unfavorable geometry of the 1138 comet affected its observing conditions in three different manners: the comet was getting fainter as the geocentric distance kept increasing, but its light was also dimmed by backscattering. In addition, as the rate of angular separation from the Sun was slowing down by the growing distance from the Earth, the comet was projecting on a relatively brighter sky background. Conversely, the favorable geometry helped the 1106 comet look brighter not only because of the decreasing geocentric distance (the first seven days), but also on account of forward scattering, especially in the early hours after perihelion. And because the comet was receding fairly rapidly from the Sun in the sky, the period of daylight appearance was cut short.
In the following, the observing conditions of the 1106 and 1138 comets are examined quantitatively. To establish whether the star-like head of a sungrazer could be detected in broad daylight with the naked eye, I employ Schaefer’s (1993, 1998) algorithm for the visual limiting magnitude of stellar objects. Its application requires that certain approximations be introduced first.
3.1. Adopted Light Curves of the 1106 and 1138 Comets
A procedure for generating approximate light curves of historical sungrazers or their candidates, introduced in Sekanina (2022), was based on a limited amount of information available on the light curves of a few bright sungrazers, mainly Ikeya-Seki and the Great September Comet of 1882, to a lesser extent also the Great March Comet of 1843. Broadly, the formula was in line with the well-known fact that the Kreutz sungrazers were fading rapidly after perihelion unless they split at perihelion: the rate of fading was assumed to correlate with the number of persisting nuclear fragments seen after perihelion.
Applied below is a slightly expanded version, which includes a phase correction. At a zero phase angle, when the correction by definition vanishes, the light curves of the 1106 and 1138 comets were assumed to comply with two postulates: (i) they were identical before perihelion, predicated on the premise that the nuclei of the 1843 and 1882 sungrazers comprised, respectively, about 40 percent of each parent’s mass; and (ii) the 1138 comet, a Population II member, split at perihelion into more persisting fragments than the 1106 comet, a Population I member. Because of the fragmentation, the light curves were not, in general, symmetric relative to perihelion.
Given that the absolute preperihelion magnitudes of the Great March Comet of 1843 and the Great September Comet of 1882, normalized to 1 AU from both the Sun and Earth, $H_{0}^{-}$, were adopted in Sekanina (2022) to equal 3.5 and 3.4, respectively, the absolute preperihelion brightness of the parents should be greater in proportion to the surface area, or mass to the power of $\frac{2}{3}$, that is $(1/0.4)^{\frac{2}{3}}\!=\!1.84$, or by 0.66 mag. Averaging, I get for both the 1106 and 1138 comets $H_{0}^{-}\!=3.45-0.66\simeq 2.8$. In line with the observed light curve of Ikeya-Seki, the preperihelion brightness variations with heliocentric distance $r$ were assumed to follow an $r^{-n^{-}\!}$ law, where $n^{-}\!=4$. The apparent preperihelion magnitude, $H_{\rm app}^{-}(r,\Delta,\alpha)$, for either of the two 12th century comets was given by the expression
$H_{\rm app}^{-}(r,\Delta,\alpha)=2.8+10\log r+5\log\Delta+\Phi(\alpha),$ (1)
where $\Delta$ is the geocentric distance, $\alpha$ is the phase angle, and $\Phi(\alpha)$ is the phase correction, computed from the standard law for dust-rich comets, formulated by Marcus (2007). Both $r$ and $\Delta$ are in AU and $\Phi(0^{\circ})=0$. It certainly is possible that the light curve flattens and the law in Equation (1) overestimates the actual brightness at extremely small heliocentric distances. However, sungrazers spend very little time, always less than one day, at close proximity of the Sun, so that this approximation is usually not a problem in practice.
After perihelion the intrinsic brightness is assumed to vary with heliocentric distance as $r^{-n^{+}}$, where in general $n^{+}\!\neq n^{-}$. The post-perihelion light curve must connect with the preperihelion light curve at perihelion, and the condition for the absolute post-perihelion magnitude, $H_{0}^{+}$, reads
$H_{0}^{+}\!=H_{0}^{-}\!+(10-2.5\,n^{+})\log q,$ (2)
where $q$ is the perihelion distance in AU. Along the post-perihelion branch of the orbit, the slope $n^{+}$ of the light curve is assumed to be a function of nuclear fragmentation at perihelion, decreasing with the increasing number $\nu_{\rm frg}$ of persisting fragments. A linear law, fitting the data on (a) comet Ikeya-Seki, which displayed two nuclei after perihelion and whose $n^{+}\!=4.0$, and (b) the 1882 sungrazer with five to six fragments and $n^{+}\!=3.3$ (Sekanina 2002), is given by
$n^{+}\!=4.4-0.2\,\nu_{\rm frg}.$ (3)
For objects that do not split at perihelion, such as the 1843 comet, $\nu_{\rm frg}=1$, and $n^{+}\!=4.2$. This is in line with bright sungrazers’ fading more rapid than their brightening in the absence of perihelion splitting. However, this procedure is inapplicable to the dwarf sungrazers, such as those seen in images taken with the coronagraphs on board the Solar and Heliospheric Observatory.
The probable perihelion-fragmentation histories of the 1106 and 1138 comets were addressed in one of my recent papers (Sekanina 2025). I speculated that besides the Great March Comet of 1843, the main returning mass of the Great Comet of 1106, comet C/1668 E1 was a major preceding fragment and another large piece could arrive at perihelion around the year 2050. Their possible early sibling — a comet of 1529, and another one, which might arrive in the mid-23rd century, are just extremely remote possibilities. Adopting $\nu_{\rm frg}=3$, I get $9.52.5\,n^{+}\!=$ 9.5 from Equation (3). With the perihelion distance from Table 1, the absolute post-perihelion magnitude of the 1106 comet is $H_{0}^{+}\!=+1.7$ from Equation (2).
In the same recent paper I proposed that besides the Great September Comet of 1882 and comet Ikeya-Seki, another likely fragment of the Chinese comet of 1138 was X/1702 D1 and one to return to perihelion before 2040, possibly even before the end of this decade. Uncertain as a related piece was a “sun-comet” recorded in China’s Shandong Province in April or May, 1792 (Strom 2002). I adopt $\nu_{\rm frg}=4+$ and $2.5\,n^{+}\!\simeq 8.9$. For the absolute magnitude I get $H_{0}^{+}\!=+0.5$. Accordingly, the post-perihelion light curves become
[…]
The assumptions of an identical preperihelion light curve with the 1106 comet and a stronger propensity for fragmentation at perihelion leads to the 1138 comet being intrinsically brighter after perihelion, by more than 1 magnitude at 1 AU from the Sun.
Figure 1.— Conditions for daylight observations of the 1106 and 1138 comets in the first 10 days after perihelion. Even though the 1138 comet was intrinsically brighter after perihelion, the greatly superior observing conditions made the 1106 comet substantially brighter in daylight, as shown by comparing the light curves (thick lines) with the limiting magnitudes for the naked eye (thin lines) at local noon. The shaded area predicts that the favorable daylight conditions for detecting the 1106 comet may have prevailed over the first three days after perihelion. The 1138 comet is predicted to have been undetectable with the naked eye in daylight.
3.2. Site Data For Limiting-Magnitude Prediction
In Schaefer’s algorithm, the visual limiting magnitude is a function of four parameters that depend on the observing site and the conditions at the time. For the exercise I choose the only location where both comets were observed, which was the territory of the Sung Dynasty in China. The geographic data for the putative observing site should adequately be approximated by an average for two modern Chinese cities: Nanjing, the capital of Jiangsu Province, and Hangzhou, the capital of Zhejiang Province.
The data needed for the algorithm are the geographic latitude, for which I use +31; the altitude above sea level of 20 meters. For daylight conditions I adopt the air temperature and relative humidity of, respectively, 9C and 74% in early February for the 1106 comet and 31C and 79% in early August for the 1138 comet.
3.3. Visibility of the Comets in Broad Daylight
The effects of the Moon in broad daylight were ignored, because when near the comet in the sky, the Moon must have been near new and its contribution negligible. The Moon was new on 1106 February 5.95 UT and on 1138 August 7.62 and September 6.29 UT. Except near the horizon, the limiting magnitude depends only weakly on the Sun’s and comet’s zenith distances, and I considered the case of both the Sun and the comet near the local meridian and their zenith distances equal. For such conditions, the implied zenith distances were $\sim$50 for the 1106 comet and $\sim$10 for the 1138 comet.
The apparent light curves of the two comets and the respective limiting magnitudes are plotted in Figure 1. The striking feature is the reversed order of the light curves. Whereas Equations (4) showed the 1138 comet to be more than 1 mag intrinsically brighter after perihelion, Figure 1 predicts that the 1106 comet should have appeared distinctly brighter to the terrestrial observer. The driver of this reversal was the differential phase effect early after perihelion, while later the disparity in the geocentric distance became more important. And, as follows from Table 2, for a given heliocentric distance (which governs the intrinsic-brightness variations) the solar elongation of the 1138 comet was always only about one half of the solar elongation of the 1106 comet, a major factor affecting the visibility of the two objects.
An important point to make is that the unfavorable geometry appears to affect the visibility of a sungrazer most severely in close proximity of the Sun. Indeed, while the 1106 comet was brighter than the limit for naked-eye detection over the first three days after perihelion, the 1138 comet stayed in daylight below this threshold at all times. The experience with the 1138 comet, including its comparison with the 1106 comet, thus suggests that the widely-accepted perception that a sungrazer at perihelion between mid-May and mid-August will be missed, unless it can be seen in daylight is misleading and should be replaced with will be missed in daylight and the rest of the statement, namely, will never be seen in a dark sky should be deleted. These comments fundamentally change the status of the sungrazers passing through perihelion between mid-May and mid-August.
Figure 2.— Projection onto the plane of the sky of the predicted orbit for X/1106 C1 close to perihelion (equinox J2000). The dotted part of the orbit indicates that the comet was then behind the Sun. The bullets show the locations of the comet at the times, reckoned in hours from perihelion. Parenthesized are the respective phase angles. Also shown are arcs of solar elongations of ${}^{\circ}\!\!$.6, ${}^{\circ}\!$, and ${}^{\circ}\!\!$.6. Receding from the Sun at a nearly constant projected rate of $\sim$4 per day starting shortly after perihelion, the comet was observed at Gembloux as a bright star over a period of six hours.
4. Daytime Observation of X/1106 C1
at Gembloux, Belgium
The extremely favorable daylight observing conditions predicted for the 1106 comet appear to be confirmed by the historic record, written by Sigebert de Gembloux in his Chronicon sive Chronographia in AD $1$. A monk in the Benedictine Abbey of Gembloux, Belgium, he noted in the chronicle that on 1106 February 2 a star appeared during the daytime, between the third and ninth hour, about a cubit from the Sun. [^2] To answer the question of what was the distance in degrees is not straightforward. One can find a lot of information on the cubit in the literature, including the differences between the Sumerian cubit, Egyptian royal cubit, Greek cubit, Roman cubit, etc. and anthropologists still write scientific papers on the cubit in the 21st century (e.g., Stone 2014). Yet, all this information is irrelevant to Sigebert’s observation, because the primary usage of the term cubit in the antiquity was as a unit of length, equal to the extent of the forearm from the tip of the middle finger to the elbow. Different cubits had different lengths, mostly between 17.5 and 21 inches, or 44 and 53 cm. — Obviously, the cubit as a measure of length and the cubit as a measure of angular separation are two very different quantities, and it is unfortunate that for some strange reason the same term is used for both of them. It is more difficult to find information about the cubit measuring angular separation. And this information is not only scarce but, worse, also inconsistent. When trying to google the term, one usually finds that 1 cubit is approximately 1, with no details or reference being given. There are, however, exceptions. To an inputed request “cubit angular measure” Google’s AI Overview returns: A “cubit” as an angular measure roughly corresponded to a finger’s width held at arm’s-length, which subtends an angle of about 2 degrees. Similarly, in Wikipedia’s article on “Cubit” one can read: There is some evidence that cubits were used to measure angular separation. The Babylonian Astronomical Diary for 568–567 BCE refers to Jupiter being one cubit behind the elbow of Sagittarius. One cubit measures about 2 degrees, with a reference to Steele (2008), noting that the relationship between the cubit as a unit of length and a unit of angular separation was not in the book elaborated on. Interestingly, I was unable to find the cubit, as an angular measure, in Wikipedia’s article “List of obsolete units of measurement,” or among the angular measure units in Wikipedia’s article “List of unusual units of measurement,” or in any of the associated articles “Ancient Arabic (Egyptian, Greek, Mesopotamian, Roman) units of measurement.” (I have not investigated sources concerning the works by Otto E. Neugebauer, as this would involve prohibitive time demands.) On the other hand, in their paper on ancient Babylonian time measurements Stephenson & Fatoohi (2008) wrote that it should be noted that observed angular distances between celestial bodies were almost invariably expressed by the Babylonians in terms of the kùš (cubit), which was equivalent to between about 2 and 2.5 degrees. The various sources clearly provide a rather wide range of numbers. Using the defining ratio of the finger’s width to arm’s-length, taking the standard values of 1.85 cm for the finger’s width and $\frac{14}{9}$ for the ratio of arm’s-length to forearm’s-length of 46 cm, I derive that 1 cubit = ${}^{\circ}\!$.5, yet another estimate.
Although the comet was called a star by Sigebert, a tail up to 100 long several days later left no one doubting the nature of the object (Section 5). It is common that in broad daylight only a very short or no tail is seen with the naked eye because of the low surface brightness. What is remarkable about the statement is the time constraint. Sigebert clearly reckoned the time from sunrise. He said nothing about the first three hours, but that means nothing; the Sun may have been shrouded in clouds or fog, or the separation was too small for optical discrimination in the Sun’s glare, or it simply did not occur to him (and/or others) to look up the sky earlier.
In any case, to inspect the timeline more closely, I plot in Figure 2 the orbit of the 1106 comet in projection onto the plane of the sky from 3.5 hours before perihelion to nearly 13 hours after perihelion. For approximately one hour, the comet was hidden behind the Sun, including its passage through perihelion. When observed at the Abbey by Sigebert himself and/or his colleagues, the comet was already past perihelion and moving to the southeast of the Sun at a nearly constant rate of $\sim$10 per hour in the general direction of the Earth.
It turns out that from Sigebert’s description one can approximately establish the UT times of observation and therefore the comet’s perihelion time. Given that the geographic longitude and latitude of the Abbey were, respectively, ${}^{\circ}\!$.70 E and ${}^{\circ}\!$.56 N, and the elevation was 150 meters above sea level, the Sun rose at 7:13 UT and set at 16:42 UT on 1106 February 2 (not corrected for refraction). Daylight thus extended over 9:29 hours and, rounded to the nearest quarter of an hour, the comet was seen between 10:15 and 16:15 UT. Since the latter time was less than 30 minutes before sunset, it is reasonable to conclude that the event’s reported termination was a corollary to the deteriorating atmospheric conditions at low elevations, not the comet’s fading. During the six-hour observing window, the comet’s solar elongation, $\epsilon$, is predicted from Figure 2 to have increased by slightly more than 1. More precisely, the relationship between $\epsilon$ and the time from perihelion, $\tau$ (in days), is given by an empirical formula:
$\epsilon$ $=$ $-0^{\circ}\!.1904+4^{\circ}\!.3822\,\tau-0^{\circ}\!.3351\>\!\tau^{2}\!,$
[…]
with a mean error of ${}^{\circ}\!$.0044 or ${}^{\>\!\prime\!}$.26 and applicable to an interval of +0.18 day $\leq\!\tau\!\leq$ +0.80 day. For example, if 1 cubit is taken to equal ${}^{\circ}\!$.1, the comet’s solar elongation would have been ${}^{\circ}\!$.6 some 4.4 hours after perihelion and ${}^{\circ}\!$.6 about 5.7 hours later. The elongation would equal 1 cubit 7:15 hours after perihelion, equivalent to the middle of Sigebert’s six-hour long window, 13:15 UT. One finds in this case that the comet passed perihelion at 6:00 UT, or on February 2.25 UT.
Similarly, if 1 cubit equaled 2, the comet would have reached the solar elongation of 1 cubit 12:30 hours after perihelion on February 2.03 UT, about 5.5 hours earlier than in the previous case. To summarize, the uncertainties in the source data (Sigebert’s rounding off the times to 1 hour; giving the solar elongation as about 1 cubit; the large error in the cubit-to-degree conversion; etc.) suggest that the perihelion time has been established with estimated accuracy of $\pm$0.3 day.
Although it was rather obvious from Figure 1 that just hours after perihelion the 1106 comet’s brightness should have greatly exceeded the visibility threshold near local noon, it remained unanswered whether this also applied at larger zenith distances. Besides, Figure 1 was plotted for a site at a geographic latitude very different from that of the Belgian site. To address this problem I computed the comet’s apparent magnitude from the adopted light curve in Equation (4) as a function of time between sunrise and sunset on 1106 February 2 and then compared it with the limiting magnitude in Figure 3. As the site constants for the Abbey of Gembloux I employed the geographic coordinates and elevation mentioned above, and adopted an average temperature of +4C, [^3] The limiting magnitude depends only weakly on the temperature (with typical rates on the order of 0.01 mag per ${}^{\!\circ}$), so that a constant-temperature approximation is, except near sunrise and sunset, deemed acceptable. and relative humidity of 89%. Over the period of 9.5 hours between sunrise and sunset the comet’s brightness is predicted to have dropped by $>$6 mag, its solar elongation to have increased by almost ${}^{\circ}\!\!$.5, and its zenith distance to have stayed above 66.
Figure 3.— Light curve (thick line), limiting magnitude (thin line), and solar elongation (dotted line) predicted for comet X/1106 C1 vs time on 1106 February 2, when the comet was observed as a bright star over a period of six hours in broad daylight at close proximity of the Sun, according to a historical record by Sigebert, a monk in the Benedictine Abbey of Gembloux, Belgium. The comet’s brightness, greatly enhanced by the effect of forward scattering (up to nearly 4 mag at maximum) is predicted to be some 4–6 magnitudes above the visibility threshold for the naked eye, except at times of very low elevation above the horizon, just after sunrise and before sunset. The start and end of the six-hour long period of observation are known with only limited accuracy. The comet was assumed to pass perihelion at 6:00 UT on February 2.
Figure 3 shows that the comet’s brightness stayed above the predicted visibility threshold by 4–6 magniudes except when very low above the horizon. However, it should be noted that the comet profited greatly from effects of forward scattering, which at peak augmented its brightness by nearly 4 magnitudes, about 3 hours after perihelion, and contributed more than 2.5 magnitudes throughout the period of time between 7:30 and 16:30 UT. Interestingly, the end of Sigebert’s observation coincides with the time when the comet’s light curve meets the limiting magnitude. The inaccuracies in Sigebert’s report affect choice of the perihelion time, which influences the positions of the three curves in Figure 3. For an earlier time, the comet’s predicted solar elongation would be greater and the comet would be fainter but the detection limit would drop as well. In any case, the comet must have offered an incredible sight and Sigebert, in retrospect, left a valuable piece of scientific data.
An interesting question to which I do not have an answer is how much of a problem for this interpretation is Sigebert’s silence about the comet’s increasing solar elongation over the six-hour period. Is it reasonable to expect that the gradual increase by up to 1 was large enough that it should have been detected, or is this merely evidence that Sigebert’s observations were not very careful? Or could his silence favor a larger separation (i.e., more degrees per Sigebert’s cubit), because an increase in the separation by the same amount is then harder to notice?
5. Tails of the 1106 and 1138 Comets: Comparison
Under favorable conditions, a massive sungrazer passing perihelion does not need more than just a few days to develop a tail, which becomes spectacular over a short period of time and brighter than the comet’s head for weeks. The main contributor in terms of the light effect is dust, while plasma is present mainly early after perihelion. For the terrestrial observer to witness a prominent dust tail, three conditions must be satisfied. Any sungrazer complies with the first condition: in order to produce a long tail, microscopic dust grains that populate it must be rapidly transported to gigantic distances from the comet’s nuclues, the task accomplished by immensely high accelerations furnished by solar radiation pressure at very small heliocentric distances. If the acceleration on a submicron-sized particle equals, say, 0.4 cm s-2 at 1 AU from the Sun, it skyrockets to 40 m s-2 at 0.01 AU from the Sun. The second condition is satisfied only by massive sungrazers: sufficiently large amounts of microscopic dust must be injected into the tail to make it prominent, which a small fragment nucleus is incapable of. And the third condition requires that the tail point towards the Earth: the effect perceived by the terrestrial observer becomes then overwhelming, because the tail not only appears brighter and its projected extent more impressive on account of its smaller geocentric distance, but also because by virtue of its being located between the Sun and Earth it is subjected to effects of forward scattering, thereby further enhancing its brightness. The 1106 and 1138 comets are nearly extreme examples that demonstrate the utmost importance of the geometry.
One prerequisite for examining the dust tails is knowledge of the properties of the debris they contain. Fortunately, the highly relevant data — on the tails of the Great March Comet of 1843, the Great September Comet of 1882, and comet Ikeya-Seki of 1965 — have extensively been studied (e.g., Schwab 1883, Bennett & Venter 1966, Hagughney et al. 1967, Matyagin et al. 1968, Weinberg & Beeson 1976a, 1976b, Krishna Swamy 1978, Warner 1980, Krishan & Sivaraman 1982, Sekanina 2023). Also investigated have been the dust tails of a number of dwarf sungrazers (e.g., Sekanina 2000, Thompson 2009, 2015), imaged by the coronagraphs on board the Solar and Heliospheric Observatory (SOHO); most of these objects appear to be debris of the 1106 comet (Sekanina 2024).
The results on the sungrazers’ dust tails are remarkably consistent in that they show that the highest radiation pressure accelerations that particles are subjected to are close to $\beta=0.6$ the solar gravitational acceleration, compatible with the presence of submicron-sized silicate grains, primarily olivine and pyroxene. The other important finding is that all three major sungrazers had a mixed tail, including both dust and plasma, at least early after perihelion. The plasma tail was then extending farther than the dust tail, thus determining the feature’s overall length. As the heliocentric distance grew, the plasma tail was gradually disappearing.
[…]
Based on the constrained radiation-pressure accelerations of grains, the dust tails of the 1106 and 1138 comets are modeled schematically in Figure 4, which displays the tails six days after perihelion with an identical system of syndynames projected under the early February and early August conditions. The difference is astonishing. In February, when the tail debris was expanding through space in a general direction of the Earth, the 1106 comet managed to build a spectacular tail that extended over half of the sky. On the other hand, in August, when the motion of the 1138 comet was generally away from the Earth, the projected tail exhibited very sharply bent contours. It was pointing away from the Sun near the head, but its orientation then reversed by nearly 180, in the direction of the Sun, in the proximity of which the feature terminated. The hook-like shape was entirely a projection effect that contributed to the comet’s lackluster appearance at the time it was supposed to be at its brightest.
The enormous disparity between the properties of the projected dust tails of the 1106 and 1138 comets in the early post-perihelion period is apparent from Table 3, in which I show, as a function of observation time, the heliocentric and geocentric distances of particles subjected to radiation pressure acceleration of 0.6 the solar gravitational acceleration; their projected distance (as a tail length) and direction (as a position angle for equinox J2000) from the nucleus; and their phase angle. It should be pointed out that the ejecta released within a few tens of minutes of perihelion are subjected to strong sublimation effects and do not contribute much to the post-perihelion tail, as confirmed by the orientations of the observed tails. Comparison with the position angle of the radius vector in Table 2 shows, for example, that one day later the perihelion ejecta of the 1106 comet projected in a direction only 50 away from the sunward direction. At different times after perihelion, dust of different ejection times contributed to the tail differently, but the tail length was determined by the ejecta that left the nucleus only a fraction of the day after perihelion.
Figure 5.— The positions of comet X/1106 C1 in the sky at sunset on 1106 February 3–13 at the geographic latitude ${}^{\>\!\!\circ}$ and the longitude 120E, an approximate location of the capital of the Sung Dynasty (near today’s cities of Hangzhou and Nanjing). The Sun’s azimuth at sunset was about 75, measured from the south westward. The plotted tail arcs illustrate the feature’s orientations, not lengths. The Sun’s dimensions are strongly exaggerated.
6. Comet X/1106 C1 at and After Sunset
Continuing with the investigation of the observing conditions in twilight, I plot in Figure 5, in the horizontal coordinate system, the positions of the Great Comet at sunset between 1106 February 3 and 13. With the Chinese observations in mind, the plot refers to a longitude and latitude of, respectively, 120E and ${}^{\>\!\!\circ}$, on the assumption that the comet was at perihelion on February 2.25 TT. Over the ten days, the sunset times varied from 9:40 to 9:48 UT, the solar elongations of the comet from ${}^{\circ}\!$.5 to ${}^{\circ}\!$.3, and its elevations from ${}^{\>\!\!\circ\!}$.5 to ${}^{\circ}\!$.2.
Although the head of the comet was not bright enough to be detected at sunset, Figure 5 shows that the observing conditions at the northern latitudes near 30 were rapidly improving with time in early February. This applies not only to the Chinese observations, but also to those from Palestine (e.g., the city of Jerusalem is at nearly the same latitude) and other locations. The large recorded disparities in the tail length could be explained by the wide variety in both the sites and times of observation. Reports of shorter tail lengths could be expected from more northerly locations and from later, post-twilight observations, when the comet’s head was deep below the horizon and only distant parts of the tail could still be seen. This means that the best estimates of the tail length are the longest ones, near 100. The reports that the comet was “split into many pieces,” or “appeared like a broken-up star,” or like “a beam forking into many rays,” provide likely evidence for the presence of a plasma tail and its complex morphology. Although in the equatorial coordinate system the predicted position of the plasma tail was in the east-southeast direction (varying from position angle of 116 on February 5 to 103 on February 13), in the horizontal system at latitude ${}^{\>\!\!\circ}$ the tail should have been seen to point up and east, deviating at sunset from the zenith direction by 63 on February 5 and by 55 on February 13.
Figure 6.— Comet X/1106 C1 on 1106 February 9 over a period of 80 minutes, as seen from a location of geographic longitude 120E and latitude ${}^{\>\!\!\circ}$. Sunset is predicted to have occurred at 9:45 UT, when the comet was ${}^{\circ}\!$.1 from the Sun at position angle of 108 and elevation of ${}^{\circ}\!$.4. The comet set 67 minutes later, in the course of astronomical twilight, with the Sun 14 below the horizon.
To depict the observing conditions as the 1106 comet was approaching the horizon following the sunset, I plot in Figure 6 its apparent motion over a period of 80 minutes on the evening of February 9, predicted for a location at geographic longitude of 120E and latitude of ${}^{\>\!\!\circ}$. At the sunset, which was found to have taken place at 9:45 UT, the comet was ${}^{\circ}\!$.1 from the Sun at position angle of 108 and elevation of ${}^{\circ}\!$.4. The comet set 67 minutes later, in the course of astronomical twilight, with the Sun 14 below the horizon. Under these conditions the long tail must have offered a magnificent spectacle.
The potential visibility of the comet’s head in twilight was another issue of interest. I already noted that the head was too faint to detect at sunset. As the Sun was moving deeper below the horizon, the conditions were improving. Unfortunately, the comet’s elevation was also diminishing. One would expect that the optimum conditions may take place somewhere midway between sunset and setting of the comet, a guess that is supported by quantitative examination in Table 4. For the perihelion time of 1106 February 2.25 and the selected observing site in China, the post-sunset conditions were compared for February 9 (sunset at 9:45 UT) and February 13 (sunset at 9:48 UT). The predicted apparent magnitudes of the comet were $-2.5$ mag and $-0.8$ mag, respectively, and the effects of interference by the Moon were accounted for; it was 3.5 days old and 27 from the comet’s head on February 9, 7.5 days old and 68 away on the 13th. The table shows that the comet’s head is predicted to have been visible to the naked eye over a period of 40 minutes (from 15 to 55 minutes after sunset) on February 9 and over a period of 77 minutes (from 17 to 94 minutes after sunset) on February 13. At the most favorable times on the two days the comet’s head was found to be, respectively, a little over 2 mags and 3 mags above the naked-eye visibility threshold.
In summary, the present examination suggests that the sungrazing orbit for the Great Comet of 1106 offered favorable conditions for its observing both in daylight at close proximity to the Sun in the first three days following the perihelion passage and under twilight conditions on subsequent days, when sightings of the comet’s head were constrained to fairly short intervals of time but the long, brilliant tail — apparently consisting of both the plasma and dust constituents — was what made the comet great. The favorable geometry contributed significantly: the comet approached the Earth to 0.8 AU about one week after perihelion and parts of the dust tail to less than 0.5 AU a few days earlier. The small geocentric distances made the tail look larger, as well as the head and tail brighter, in the earliest post-perihelion days also because of strong effects of forward scattering by microscopic dust. The bulk of the comet’s recorded observations, made in daylight, twilight, and nighttime, are consistent with the results of this exercise, supporting the suggestion that this object indeed was a Kreutz sungrazer.
My final point on the 1106 comet concerns the time of perihelion, for which Hasegawa & Nakano (2001) found January 26 $\pm$ 5, following their reexamination of an earlier result, January 30 $\pm$ 5, by Hasegawa (1979). The large error notwithstanding, Hasegawa & Nakano’s perihelion time could not fit the Gembloux observation, because the comet would no longer be in close proximity of the Sun on February 2. In fact, the time spent by the object, before and after perihelion, within 1–2 from the Sun is seen from Figure 2 to have been shorter than 24 hours, which essentially rules out all perihelion times before about February 2.0 TT.
7. The Comet of 1138 at and before Sunrise
Turning now to Population II, the reader will remember my conclusion from Section 3.3 that the comet of 1138 was too faint to detect in daylight shortly after perihelion. The plasma tail could not then be observed and the appearance of the dust tail certainly did not contribute anything to facilitate detection (Figure 4).
The 1138 comet was located to the west of the Sun as it was emerging from perihelion, becoming eventually a morning object. The comet’s very small solar elongations, barely 15 some 10 days after perihelion, shown in Table 2, did not offer much optimism, but the issue not addressed by the table was the comet’s position relative to the Sun in the horizontal coordinate system. Examining the comet’s location at sunrise, I now followed an avenue of inquiry similar to that employed for the 1106 comet. The results, again applicable to the selected Chinese site, turned out to be extremely disappointing (Table 5). Fully 11 days after perihelion the comet was less than 10 above the horizon when the Sun already was at the horizon. Under such conditions, there was no hope whatsoever to pick up the comet with the naked eye as a twilight object in the early post-perihelion days.
What happened then that only three weeks later, in early September, the comet did get discovered? The answer has three parts: (i) a gradually increasing solar elongation that prompted the visibility threshold to drop much more rapidly with time than was the fading rate of the comet; (ii) an intervening event increasing temporarily the tail’s surface brightness; and (iii) favorable tail-projection developments. I now comment on each point.
The best way to document the first point is to compare the observing conditions at a few relevant times. I choose the above date of August 12 and the all-important date of September 2 (UT, i.e., September 3 local time) — the day of observation. In addition, August 20 was selected as the third date, because the Earth was then crossing the comet’s orbital plane. Focusing first on the comet’s head, Table 6 shows the relation between the predicted apparent magnitude and the limiting magnitude. On August 12 the comet is predicted to have been at least 5.5 mags fainter than the visibility threshold, on August 20 at least 2.5 mags fainter. In early September, at the time of recorded observation, the comet’s brightness is predicted to have been comparable to the limiting magnitude [^4] It is noted that these numbers do differ a little from those in Sekanina & Kracht (2022) because of slightly different laws for the light curve used, now including the phase term. over a short period of time, about midway between the rise of the comet and sunrise. The comet’s head is predicted to have been at this time slightly brighter than magnitude 3, about 10 above the horizon, and 36 from the Sun, which was some 16 below the horizon. The computed limiting magnitude is of course subject to rather large uncertainties for a number of reasons, so that the head may of may not have been seen in early September. There is in fact a good chance that it was not, the argument being offered below.
The recorded sighting is that of a hui comet, that is, one with a tail. On September 2 the comet already was 1.76 AU from the Earth, but only 1.12 AU from the Sun, so it is likely that both the plasma and the dust tails were present. The syndynames in the dust tail for September 2.8525 UT (shown in italics in Table 6) are displayed in Figure 7, which suggests that, compared to August 7 (Figure 4), the tail’s projected figure remained essentially unchanged, except for its considerable expansion.
Of distinct assistance to the projected dust tail’s improved visibility was the Earth’s transit across the comet’s orbit plane on August 20, which helped increase its surface brightness; this was the intervening event mentioned above under (ii). As a result, the line of sight from the Earth made only ${}^{\circ}\!$.7 with the orbit plane at the time of observation. Also, because of the orbit plane crossing, the tail was curving in the opposite sense on September 2 (Figure 7) compared to August 7 (Figure 4).
[…]
Because all dust ejected about 0.1 day after perihelion and later is seen in Figure 7 to have by early September been projecting at larger distances from the Sun and at greater elevations than the comet’s head, it was this part of the tail that had the best chance of detection at that time. This is my argument under point (iii) above. In particular, one can see from the figure that the dust ejected near 1 day after perihelion was in fact located at an elevation about 6 higher in the sky than the head. It is hard to estimate how close to the horizon the feature could have been followed, but chances were that the visible tail length was at most 10 or so. It is also likely that the predicted marginal naked-eye detection of the head (Table 6) meant in practice that it did not stand out enough to be readily noticed. Its position, sideways from the main body of the dust tail and near the middle along its long axis, was very unusual and to the medieval observer could have been puzzling to the point that he probably would have commented on it.
The most unusual conditions of tail projection in September are further underlined by the data presented in Table 7, which likewise assist in estimating the tail length in early September and also help explain why the comet was seen over a rather limited period of time. The striking features of the table are the enormous distances from the Sun and even more so from the Earth reached by the submicron-sized dust particles (subjected to a radiation pressure acceleration of 0.6 the solar gravitation acceleration) ejected in the first hours after perihelion. (Such grains dominated the tails of the great Kreutz sungrazers, such as Ikeya-Seki, early after perihelion.) For example, the particles ejected a little less than one hour post-perihelion, which in Figure 7 make up the part of the tail appearing just barely above the horizon, were located more than 6 AU from the Sun and more than 7 AU from the Earth on September 2 UT and between 11 and 12 AU on September 28 UT, i.e., the morning hours of September 29 local time, when the comet went out of sight (Ho 1962). On the other hand, the heliocentric and geocentric distances of the microscopic dust ejected as late as 24 hours after perihelion, which occupies the area of the tail near its highest point in Figure 7, were only between 2 and 2.7 AU on September 2 and between 3.6 and 3.9 AU on September 28; of course, the production rate of dust was much higher one hour after perihelion than 23 hours later.
The presence of the plasma tail should have further facilitated sightings of the comet in early September. If the plasma tail extended essentially along the radius vector, the relation between its angular length $\Theta$ and linear length $\Lambda$ was a function of the comet’s geocentric distance $\Delta$ and the phase angle $\alpha$:
$\tan\Theta=\frac{\Lambda\sin\alpha}{\Delta\!+\!\Lambda\cos\alpha}.$ (6)
With $\Delta\!=\!1.76$ AU and $\alpha\!=\!32^{\circ}\!$.1 on September 2.85 UT, one gets $\Theta\!\simeq\!7^{\circ}$ for a tail 0.5 AU long and $\Theta\!\simeq\!11^{\circ}\!$.5 for a tail 1 AU long. It therefore cannot be ruled out that the plasma tail appeared a little longer than the dust tail in the antisolar direction, reaching perhaps up to nearly 5 farther above the horizon than the dust tail. By late September the heliocentric distance increased to almost 1.7 AU and the plasma tail should have been waning and with the geocentric distance approaching 2 AU, its length should have been curtailed, making its detection no longer likely.
These considerations and results of the modeling of observing conditions for the comet of 1138 lead to two important conclusions that revise the currently recognized notions about the Kreutz sungrazers reaching perihelion at the “wrong” time of the year (Section 2). It is likely that the conclusions apply to the entire period from mid-March until mid-August, even though they strictly refer only to the weeks around August 1.
The first conclusion covers the problem of detection. A Kreutz sungrazer reaching perihelion at the “wrong” time of the year cannot be discovered with the naked eye in broad daylight near the Sun, at its brightest, because it is not bright enough. On the other hand, there appears to be a narrow window of opportunity to see such a comet with the naked eye even at low to moderate northern latitudes starting about one month after perihelion, as a morning object low above the eastern horizon. The optimum conditions are strongly constrained not only in terms of the observing period (a few weeks), [^5] The severity of this limitation depends on the fading rate of the comet, for the northern-hemisphere observers it is also affected by the comet’s modest southerly motion. The declination of the 1138 comet was close to ${}^{\circ}\!$.5 on September 2, but close to ${}^{\circ}\!$.5 on September 28 (equinox of the date). but especially in terms of the time of the day, an interval whose length is curbed by a range of elevations of the comet allowed between its rise on the one hand and progressing twilight prior to sunrise on the other hand.
In this context, comparison with the sungrazing comet Pereyra (C/1963 R1) is illuminating. It passed perihelion on September 23, merely three weeks later in the year than predicted for the comet of 1138. Accordingly, Pereyra’s observing conditions were more favorable, but not by much. Also, Pereyra was not a member of Population II, so that the geometry was slightly different. Yet, while intrinsically brighter than Ikeya-Seki, [^6] From sets of visual brightness observations I determined the post-perihelion absolute magnitudes of 5.9 for Ikeya-Seki and 4.4 for Pereyra (Sekanina 2002); these compare with an adopted absolute magnitude of 0.5 for the comet of 1138. Pereyra was not detected in daylight, [^7] If the arrival of comet Pereyra were delayed by 16 years (less than 2 percent of its orbital period), it would have saturated the imager on board the Solwind satellite. but was discovered three weeks after perihelion. Over a period of several days following its discovery, Pereyra was a naked-eye object, displaying a tail several degrees in length. Although the conditions were clearly more favorable in the southern hemisphere (e.g., Bennett et al. 1964), the comet and its tail were seen by northern-hemisphere observers as well. [^8] The visibility of comet Pereyra with the naked eye is apparent from the narratives by some observers [e.g., Capen (1964), who was stationed at Wrightwood, California], but also from an incredible story involving a 1977 letter written by G. E. D. Alcock, in which — as aptly recounted in 2003 by B. G.,Marsden in his lecture given in Alcock’s memory to a RAS-BAA Pro-Am Discussion Meeting: Meteorites, Meteors, & Comets and documented by Shanklin (2004) — he wondered, in retrospect, whether a thin pencil-like beam of light that he had seen low in the sky on 1963 September 12, two days before comet Pereyra was discovered, happened to be the comet’s tail. Telescopically the comet was under observation over a period of three months, with the last observation in mid-December made from the northern hemisphere. Thus, notwithstanding the various differences (such as orbital incompatibility, a gap of three weeks in the perihelion time, an intrinsic-brightness gap of almost 4 magnitudes, etc.), one notices very obvious similarities between the predicted traits of the 1138 comet and the observed traits of comet Pereyra.
The second conclusion concerns a rule of thumb to investigate a correlation between the time of last naked-eye observation and the comet’s apparent magnitude. Sightings of a bright sungrazing comet in the ancient or medieval times were those of its tail, which was seen even after the comet’s head was no longer visible to the unaided eye. I tested this relationship on the well-observed sungrazers, such as the Great March Comet of 1843, the Great September Comet of 1882, Ikeya-Seki, etc., and suggested that, based on the experience with these objects, the post-perihelion tail of a sungrazer was typically seen with the naked eye until the brightness faded to apparent magnitude of $\sim$7, so that the head became a telescopic object long before the last naked-eye tail detection (Sekanina 2022).
Applied to seven objects between AD 943 and 1695, listed by Hasegawa & Nakano (2001) as potential Kreutz sungrazers (all reaching perihelion under favorable observing conditions in Octover or November, similar to those for comet Ikeya-Seki in 1965), this exercise led to a revision of their absolute brightness downward, the correction averaging more than 2 magnitudes. The outcome was their better agreement with the absolute magnitudes of the recent Kreutz sungrazers.
A very different case on that same list is the comet of 1041, which was already briefly commented on in Section 2. Hasegawa & Nakano assigned it an absolute magnitude of --1, making it 2 mags intrinsically brighter than was my estimate for the Kreutz system’s progenitor (Sekanina 2022)! There appear to be two problems with this 1041 object, the more serious one being that two different comets — either one recorded in China as well as Korea — were almost certainly involved, as suggested by both Ho (1962) and England (2002). The first of the two comets was discovered on September 1 (Hasegawa & Nakano 2001) and “went out of sight after more than 20 days” (Ho 1962). The mentioned rule-of-thumb method offered for it absolute magnitude 4, which is 5 mags fainter than proposed by Hasegawa & Nakano. The second comet appeared in November, was seen for more than 10 days, and appears to have been intrinsically fainter. Both may have been Kreutz sungrazers, perhaps even fragments of an object that split not long before 1041.
The estimated (post-perihelion) absolute magnitude 4 for the September comet of 1041 happens to be brighter than Pereyra’s by only 0.5 mag or less and too faint for the parent, especially because circumstantial evidence suggests that the 1041 comet fragmented heavily at perihelion. [^9] The barycentric orbital period of the 1041 comet upon its arrival at perihelion was 678 years, but Pereyra’s original orbital period slightly exceeded 900 years, which implied the existence of a few fragments in an orbital-period range of $\sim$450–750 years. This would reduce the comet’s rate of post-perihelion fading, thus bringing the relationship between the absolute magnitudes of the 1041 comet and Pereyra closer to expectation. If so, its brightness should have been almost comparable to the brightness of the 1106 or 1138 comets. To achieve this, the time of the last naked-eye observation in the rule of thumb should have coincided with the comet’s apparent magnitude having equaled not 7, but about 4.5. And this is the second problem with the September comet of 1041.
Returning now to the Chinese comet of 1138, I find that around September 28.9 UT, when the comet went out of sight (see Ho 1962), its apparent magnitude computed from the second formula in Equation (4) was — astonishingly — 4.5!! A second comet that defies the rule of thumb with apparent magnitude 7 and a second comet that passed perihelion in the “wrong” time zone of the year. Given that the apparent magnitude 7 applied to the sungrazers that arrived at perihelion at “favorable” times of the year, could it be that the apparent-magnitude threshold in the rule of thumb correlates with the seasonal variations in the observing conditions because the sungrazers’ activity and the surface brightness of their tails respond to the conditions in a different manner? Unfortunately, the currently available data are too scanty and inaccurate for an in-depth investigation of this problem at this time.
8. Fragmented Nucleus of the Comet of 1138
Even though I believe that the nucleus of the Great Comet of 1106 did break up [^10] It is possible that besides the Great March Comet of 1843, at least one other sungrazer — C/1668 E1 — was also a fragment of the 1106 comet (Sekanina 2025). in the course of its perihelion passage as well, only the fragmentation event experienced by the Chinese comet of 1138 is modeled and examined here in some detail. There have been two good reasons: (i) it is virtually certain that the Great September Comet of 1882 and Ikeya-Seki were products of this event (cf. Marsden 1967), possibly with other objects; and (ii) the nuclei of the 1882 comet and Ikeya-Seki were themselves subject to fragmentation at perihelion, the first of the two splitting into five to six components; a search for similarities has been desirable.
The multiple nuclei of the 1882 comet, as well as the existence of split comets whose breakup is known to be of tidal nature but occurred at close proximity of Jupiter — such as D/1993 F2 (Shoemaker-Levy) or 16P/Brooks (C/1889 N1) — suggest that (a) the formation of a chain of more than two nearly equidistant fragments, buried in an oval cloud of debris, is by no means exceptional; and (b) as a rule, the principal fragment is — unlike in nontidally split comets — located typically not at the leading end of the chain, but rather nearer its middle. The two conditions combine to become helpful for predictive purposes. In applications, the first condition serves as an impetus in a search for conceivable fragments among potential historical sungrazers, whereas — once the principal fragment has been identified (or probably identified) — the second condition is useful in assessing a timeline of the chain of fragments, with the aim to constrain the range of fragments’ orbital periods as much as possible. Of particular interest is any hint of evidence that additional fragments could be expected to arrive in the future, especially near future.
The purpose of the following exercise is to reconstruct the possible appearance of the nucleus of the 1138 comet that the medieval Chinese observer would have seen around 4:30 in the morning of September 3 local time (or September 2.85 UT), nearly 33 days after the predicted perihelion, if — hypothetically — he had a telescope of, say, $<$similar-$\sim$ 1′′ resolution available. He probably would have seen the two condensations around the nuclei of the comet of 1882 and Ikeya-Seki, [^11] There is no correlation between the mass of a fragment and the brightness of its condensation at a particular time; a more massive fragment may be enveloped by a fainter condensation than a less massive fragment because of large, irregular fluctuations in activity. plus possibly additional condensations, if the nucleus had split into more than two sizable fragments.
The modeling of the post-perihelion, split nucleus of the 1138 comet relies heavily on two postulates: (i) the Great September Comet of 1882 was the largest surviving mass of the progenitor’s Lobe II (i.e., of Population II), a fundamental premise of the contact-binary model for the Kreutz system (Sekanina 2021); and (ii) it is not a coincidence that the difference between the barycentric orbital periods of Ikeya-Seki and the Great September Comet of 1882, amounting to 83 years, nearly equals the width of the gaps between the osculating orbital periods of the neighboring nuclei of the 1882 comet, averaging about 95 years.
In the context of the conditions (a) and (b) above, the adherence to these postulates led to the conclusion that at close proxinity to perihelion the Sun’s tidal force did, in all probability, split the nucleus of the 1138 comet not into two, but at least four or perhaps five or six components (Sekanina 2025). Chronologically, such a chain of fragments could have started with X/1702 D1 and continued with the “sun-comet” of 1792 (Strom 2002), [^12] Because of very vague nature of the “sun-comets,” this obviously is a weak entry; it is possible that the fragment expected to arrive at about this time was missed. followed by the 1882 sungrazer (the principal, most massive fragment) and by Ikeya-Seki. In order that the principal fragment be located near the middle of the chain, in line with condition (b), one can argue that at least one additional fragment should exist beyond Ikeya-Seki, and it should arrive in the near future. There is a good chance that this could happen — with the uncertainty of at least several years — around 2027, while the next one (if any) should be a matter of distant future.
The appearance of the nucleus region of a split sungrazer is essentially determined by the angular separations between the major, kilometer-sized and larger fragments, which are buried in an encompassing cloud of debris made up of smaller fragments down to microscopic dust. To describe the feature’s overall dimensions, one needs to establish the osculating orbits of the major fragments. Since a tidal breakup that involves no significant momentum exchange among the major fragments has practically no effect on their perihelion times and angular elements, they remain equal to those of the parent at the time of breakup. For the 1138 comet they are listed in Table 1. The remaining two elements vary from fragment to fragment, but the perihelion distances marginally, by a fraction of the nucleus dimensions. The only meaningful differences are in the fragments’ eccentricities, $e$, or alternatively in their semimajor axes, $a$.
The starting point in the process of modeling the appearance of the 1138 comet’s fragmented nucleus is the relationship between its osculating and future orbital periods as the previous appearance of the 1882 sungrazer, its principal fragment. From Table 1 the relevant osculating period is 754.4 yr, while the future barycentric orbital period is 1882.7 – 1138.6 = 744.1 yr. [^13] The actual calculation of the period has to be done in Julian dates (not in calendar days), which are then converted to years. This is necessary because the perihelion times in 1138 and 1882 are given, respectively, in the Julian and Gregorian calendars. When rounded off to a tenth of the year, the difference does not show up. In terms of $1/a$, the effect of the planetary perturbations of the 1882 sungrazer’s motion is
[…]
$=$ $0.012178-0.012067$
$=$ $+0.000111\;{\rm AU}^{-1}.$
It is well known that Kreutz sungrazers have no close encounters with the planets and that their motions are affected only by the indirect planetary perturbations. Because the individual fragments were close enough to one another over the post-perihelion period of time of significant indirect perturbations, I assume that their motions were perturbed equally, so that Equation (7) applies not only to the 1882 sungrazer, but to the other fragments as well. This of course is not true for their motions at times just before their next arrival, but by then the perihelion time has essentially been fixed. The prime driver of the $(1/a)_{\rm fut}\!-\!(1/a)_{\rm osc}$ difference is the heliocentric distance at aphelion, which is substantially affected by the post-perihelion, not preperihelion, perturbations.
Figure 8.— Modeled appearance of the fragmented nucleus of the 1138 comet on 1138 September 2.85 UT, nearly 33 days after perihelion. Plotted are the astrometric positions (equinox J2000) of the proposed major fragments, besides the Great September Comet of 1882 and Ikeya-Seki also the assumed fragments X/1702 D1, and the 1792 sun-comet; and a predicted sungrazer of 2027. The extent of the chain amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3 and is enveloped in an elongated cloud of dust debris. All five fragments are assumed to be active at the time to a degree that their condensations are not overwhelmed by the diffused light from the debris, an ideal case discussed below.
[…]
The last remaining step is to compute the astrometric positions of the five fragments for the time of interest. Displayed in Figure 8 is the modeled appearance of the nucleus region of the 1138 comet on September 3 around 4:30 local time (September 2.85 UT). The extent of the fragmented nucleus (from X/1702 D1 to the predicted comet of 2027) amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3. Terminating the chain of fragments instead with Ikeya-Seki, the length would at the time be ${}^{\prime\prime}\!$.2, with Ikeya-Seki and the 1882 sungrazer only ${}^{\prime\prime}\!$.8 apart. At the geocentric distance of 1.76 AU this was equivalent to merely 3600 km, but because the line of sight made with the line connecting the two objects an angle of merely 14, they were in space 14,500 km apart.
It is illuminating to compare the modeled appearance of the 1138 comet in Figure 8, at a heliocentric distance of 1.12 AU, with the observed appearance of the fragmented nucleus of the 1882 sungrazer at similar post-perihelion distances. Since it will never be known how the nucleus of the 1138 comet actually looked like on September 2 UT, this exercise is purely academic. Yet, inspection of the range of possible appearances offers an insight into the potential subsequent evolution and ultimate fate of the fragments.
Table 8 compares the two comets in terms of separation distances among the individual nuclear fragments. For the 1882 sungrazer, Kreutz (1891) was able to determine separate sets of orbital elements for four of its nuclear components by linking the normal places of the preperihelion nucleus with those of the respective post-perihelion fragment. The four condensations, designated A, B, C, and D in the standard nomenclature, were in the notation used by Kreutz referred to as points No. 1, No. 2, No. 3, and No. 4, respectively. [^14] Marsden & Williams’ (2008) catalogue lists Kreutz’s (1891) sets of nonrelativistic orbital elements for nuclei A, C, and D of the 1882 sungrazer, but Hufnagel’s (1919) set of relativistic elements for nucleus B. For the sake of uniformity, I use the sets of nonrelativistic elements by Kreutz for all four nuclei.
[…]
Figure 9.— Drawing of the fragmented nucleus of the 1882 sungrazer, made by Elkin on 1882 October 13.1 UT, as seen with a 15-cm Grubb equatoreal. This was the only time that he was able to resolve four condensations in the chain; shown are the separations that he reported (Gill et al. 1911). Kreutz (1888) assigned these positions to nuclei B through E, but the easternmost fragment was near the very end of the extended cloud of debris, and it had to be A. Note that most condensations are elongated. The line crossing the chain was marked by Elkin to show the point that was measured for position. The principal fragment was B, whose orbital period essentially equaled that of the nucleus before fragmentation.
The four data sets for the 1882 sungrazer in Table 8 refer to a selected set of the astrometric and/or physical observations by W. L. Elkin at the Cape Observatory (Gill et al. 1911), which were accompanied by dot renditions of the nuclear chain immersed in a cloud of debris. They were drawn to scale.
The first of these reproductions, in Figure 9, shows the fragmented nucleus on 1882 October 13 UT, 25.4 days after perihelion, as seen by Elgin with a 15-cm Grubb equatoreal. The extent of about 20′′ exceeds by 5′′ the length of the modeled chain of fragments of the 1138 comet. As Table 8 explains, this apparent imbalance happens to be the effect partly of a larger geocentric distance of the 1138 object, partly of its highly unfavorable projection, as the chain of fragments made an angle of only 14 with the line of sight. In reality, the 1882 fragmented nucleus was 38,000 km long at the time, less than one half the modeled 1138 nucleus.
[…]
Figure 10.— Drawing of the fragmented nucleus of the 1882 sungrazer, made by Elkin on 1882 October 22.1 UT, as seen with a heliometer. Compared to October 13, condensation A brightened significantly. The positions of the fragments B, C, and D are now those calculated from Kreutz’s orbital elements and tied to A. The overall length of the cloud of debris is more than twice as long as the distance that separates the condensations A and D. Additional, smaller, and gradually disintegrating fragments (E, F, …) must have been located beyond D. The line crossing the chain was marked by Elkin to show the point measured for position.
The observation of October 13 was the only one, when Elkin was able to resolve the nuclear chain into four fairly distinct condensations, so that the separations among all of them could be measured. Note however that most condensations appear to be elongated. From the east to the west, the measured distances were ${}^{\prime\prime}\!$.3 between the first and second condensations, ${}^{\prime\prime}\!$.7 between the second and third, and ${}^{\prime\prime}\!$.5 between the third and fourth. Kreutz (1888), who only saw the report (the drawing was published in 1911), assigned the condensations to nuclei B, C, D, and E (that is, Nos. 2, 3, 4, and 5 in his notation). However, Figure 9 clearly shows that the easternmost condensation was near the very end of the cloud of debris and must have been nucleus A.
This conclusion is supported by Kreutz’s own orbital computations, which for the three separation distances give, respectively, ${}^{\prime\prime}\!$.8, ${}^{\prime\prime}\!$.2, and ${}^{\prime\prime}\!$.6 (Table 8). The distances BA and BC agree with Elkin’s measurements to $\pm$0′′.5 and only CD is off by nearly 2′′, possibly because of a contribution to D from E. Another strong argument comes from Elkin’s rather extensive description of the easternmost (or following [^15] In the era of visual observation it was common that an object to the west of a point of reference be called a preceding object, one to the east of it a following object, apparently because when the telescope clock was stopped, the Earth’s rotation moved the star field westward.) condensation. Elkin conveys that “till the end of October [this fragment] was a fair object for observation.” However, by November 11 it became “so covered with haze, that it was obvious that soon nothing but the centre of the great mass would be available for a reference point …, and after Nov. 18 … the place of [the following condensation] was more guessed at than actually seen.” This is a consummate account of a protracted flare-up of a fragment that was undergoing a process of continuous loss of mass and should have problems to survive to the next perihelion. Certainly not the behavior of the principal nucleus B!
To document Elkin’s account, I reproduce additional three of his many drawings in Figures 10–12, which cover the period of October 22 through November 21 UT. They are annotated to indicate the relative positions of the nuclei from Table 8, as calculated from Kreutz’s (1891) orbital elements and tied to condensation A (or to what remained of it on November 21). Unfortunately, Elkin made these observations with a heliometer, whose aperture was much smaller than that of the equatoreal; accordingly, their quality is inferior.
[…]
Figure 11.— Drawing of the fragmented nucleus of the 1882 sungrazer, made by Elkin on 1882 November 1.1 UT, as seen with a heliometer. Compared to the sketch from October 22 (Figure 10), the changes were relatively subtle, except for an obvious elongation of condensation A. While C temporarily disappeared, B may have now shown up. Its computed position would be in good agreement with observation, if the center of mass of condensation A was moved a little closer to its eastern boundary than shown by Elkin. Also apparent is another condensation about 10′′ to the west of the predicted position of D, close to where one would expect condensation E (for which no orbit is available). The length of the cloud of debris increased by 10′′ in ten days.
Nonetheless, Figure 10, displaying Elkin’s drawing from October 22, does show major changes that occurred in the appearance of the nucleus over the nine days. The most obvious is the dramatic brightening and expansion of condensation A, suggesting a high rate of its mass loss, as remarked above. Another major difference is the considerable increase in the length of the feature, which, including the cloud of debris, amounted to 68′′, nearly $\frac{1}{2}$ times the distance between A and D, calculated from Kreutz’s (1891) orbital elements. A slight brightening is seen at the calculated position of condensation C, but there are no signs of B or D. As is the case with so many other split comets, no general relationship exists between fragment size and activity.
Ten days later, the appearance of the comet’s chain of fragments did not look radically different (Figure 11). The only easily perceived change was a significant elongation of condensation A, which could be explained by the presence of microscopic dust in its copious ejecta, subjected to high solar radiation pressure accelerations. If so, the center of mass of this condensation should be near its eastern boundary, and the fainter condensation several seconds of arc to the west of it would be in very good agreement with the position of the principal fragment B, derived from Kreutz’s orbital elements. And while there are no signs of fragments C or D, at least one additional condensation is apparent near the western end of the chain, at a location where one would expect fragment E. In the ten days, the cloud of enveloping debris increased its extent from 68′′ to 78′′ and was still more than twice as long as the computed distance between the fragments A and D.
[…]
Figure 12.— Drawing of the fragmented nucleus of the 1882 sungrazer, made by Elkin on 1882 November 21.0 UT, as seen with a heliometer. Major morphological changes occurred in the three weeks between November 1 and 21. Condensation A faded to the extent that Elkin was compelled to abandon using it as the point of reference for positional measurement and employ the middle of the fragmented nucleus instead. This drawing illustrates what Elkin meant by remarking that after November 18 the position of condensation A was “more guessed at than actually seen.” There are no obvious features at the computed places of fragments B, C, or D.
The final of Elkin’s four selected drawings of the fragmented nucleus was made on November 21 UT and is presented in Figure 12. Major differences in the feature’s morphology over the three November weeks are readily apparent. Especially striking is the fading of condensation A, which prompted Elkin to complain about guessing its place and moving the point of reference for his astrometric measurement to the middle of the fragmented nucleus. The extent of the cloud of debris ballooned to close to 100′′, while the computed length of the chain of fragments A–D amounted to some 53′′. No measurable features are seen at the computed locations of the fragments B, C, or D.
This cursory account of Elkin’s selected drawings of the fragmented nucleus of the 1882 sungrazer illustrates the complex evolution of, and temporal variations in, its appearance. In the context of the proposed model for the fragmented nucleus of the Chinese comet of 1138, whose principal component the 1882 sungrazer is believed to have been, unpredictable chaotic activity is not only inevitable, but in one sense downright momentous for the individual fragments’ future.
I refer of course to a chance of each fragment’s survival beyond its next perihelion as a sizable, essentially intact body. Regardless of the model used for the Kreutz sungrazer system, the 1138 and 1882 fragments cannot be readily compared because of the difference of one generation between them. In practice, this means that survival throughout the entire revolution about the Sun a priori is more likely for the former. Still, the failure to survive is not ruled out for either of them. This uncertainty is a major complication for predictions of future appearances of Kreutz sungrazers.
What helps is that we know that, in the least, two fragments of the 1138 comet did survive their following perihelion passage. And if X/1702 D1 was its fragment as well, as I assume, a total of at least three fragments survived, because the 1702 sungrazer was observed after perihelion. Further surviving fragments may arrive in the future (Sekanina 2025).
On the other hand, it is anybody’s guess, how many (and which) fragments of the 1882 sungrazer and Ikeya-Seki are going to survive their next perihelion. The principal fragments of both comets and the former’s fragment C (which with B was observed long into 1883) have the best chance, while fragment A of the 1882 sungrazer, with its highly variable activity, is an excellent candidate for early disintegration.
9. Concluding Remarks
The purpose of this paper has been to investigate the generation of Kreutz sungrazers immediately prior to the current one, that is, the direct parents. In the context of the contact-binary model (Sekanina 2021), there may exist a number of sungrazers of that generation (reaching perihelion between, say, the 10th and 13th centuries), but the foremost among them are the members of Populations I and II, the Great Comet of 1106 and the Chinese comet of 1138 (No. 403 in Ho’s catalogue), respectively. The contact-binary model was initially based on a postulate that the Great March Comet of 1843 was the largest surviving mass of the 1106 comet, whereas the Great September Comet of 1882 was the largest surviving mass of another comet, because at the time the model was conceived (late summer 2021), the 1106 comet still was the only feasible candidate for a Kreutz sungrazer of its generation. More recently, the problem of a missing second sungrazer was solved, resulting in the “discovery” of the 1138 comet and offering a virtual proof, by orbit integration, of its feasibility as a member of Population II of its generation and the parent to the 1882 sungrazer and Ikeya-Seki (Sekanina & Kracht 2022).
As the best estimate for its perihelion time is August 1, the 1138 comet can be used to examine the standard narrative that (in the era prior to Sun-exploring spacecraft) a Kreutz sungrazer reaching perihelion at the “wrong” time of the year — between mid-May and mid-August — was missed if not seen in daylight. My review of this scenario, with the comet approaching and leaving perihelion from behind the Sun, has led to a conclusion that the geometry involved dictated that it was at close proximity to perihelion — in daylight — when the singularly unfavorable observing conditions took place. This included a protracted period of extremely small solar elongations, as well as a very short dust tail, hook-shaped and turning sunward. On the other hand, as the comet receded a little beyond 1 AU from the Sun, its solar elongation increased enough and its fading was modest enough, that over a period of at most a few weeks it reached, in astronomical twilight before sunrise, elevations near 10, when sighted from a moderate northern latitude. Its dust tail, still hook shaped, stretched antisunward to elevations of $\sim$15–20, and the plasma tail may have extended up to $\sim$30 above the horizon. Accordingly, the notion that sungrazers with perihelia at the “wrong” time of the year were missed unless seen in daylight is misleading.
This was an issue with other sungrazers or potential sungrazers as well, such as the controversial comet(s) of 1041 and, perhaps surprisingly, comet Pereyra of 1963, whose perihelion passage was just outside the zone of “wrong” time. Even though intrinsically brighter than Ikeya-Seki, Pereyra was undetected in daylight and discovered only about three weeks after perihelion. The comet of 1138 thus was no exception.
The mediocrity of the daylight observing conditions for the 1138 comet is particularly well illustrated in juxtaposition with the greatly superior conditions for the Great Comet of 1106. This is not surprising, as the Earth was at almost exactly opposite points of its orbit at the times the two objects reached perihelion. Comparison has been carried out by applying a criterion based on Schaefer’s (1993, 1998) naked-eye limiting-magnitude model, leading to a conclusion that the vast differences were triggered by three causes: a disparity in geocentric distance; a related disparity in solar elongation, the 1138 comet always projecting closer to the Sun than the 1106 comet at the same heliocentric distance; and a disparity in phase correction, namely strong forward scattering by microscopic dust early after the 1106 comet’s perihelion.
In the light of these findings, the account by Sigebert de Gembloux of what turned out to be the daylight discovery of the 1106 comet on February 2 has been most welcome. The detection is not only in line with the predicted generally favorable observing conditions for this object, but even more so in the earliest post-perihelion times. The comet had apparently passed perihelion only hours before the reported observation was made.
Next I have addressed the relationship between the last glimpse of a historical sungrazer (i.e., its tail) and a predicted apparent magnitude of its head at the time. Hasegawa & Nakano (2001) assumed magnitude 5, but most of their Kreutz candidates were intrinsically much too bright. Based on reports for recent sungrazers, I concluded that the relationship followed a rule of thumb, with the tail still marginally visible when the head was long below the naked-eye detection limit and telescopically estimated at magnitude $\sim$7 (Sekanina 2022). However, all sungrazers involved in the experiment were those with highly favorable observing conditions. On the other hand, I have now found that the disappearance of the 1138 comet coincided with the head’s predicted apparent magnitude of 4.5, or 10 times brighter, in agreement with the result for the September comet of 1041. It appears that the apparent-magnitude threshold in the rule of thumb correlates with the observing conditions, an issue that is still waiting to be resolved.
The final addressed subject has been the appearance and morphology of the nucleus after its tidal fragmentation at perihelion. From limited evidence of such events at the Sun and Jupiter, it is not unusual to see the fragmented nucleus consisting of a number of nearly equidistant condensations buried in an oval cloud of debris, with the brightest ones located generally near the middle of the chain. For the comet of 1106, nuclear fragmentation has been almost a nonissue, in part because there is no evidence for splitting of the 1843 sungrazer. To my knowledge, there is only one other comet — C/1668 E1 — that could have derived from the 1106 comet, making the 1843 sungrazer the principal fragment. A vast majority of the dwarf sungrazers, detected by coronagraphs on board the Solar and Heliospheric Observatory and other Sun-exploring spacecraft, appears to be minor fragments of the 1106 comet, but that is another matter. More to the point is the possibility that additional fragments could arrive later in the 21st century and beyond.
Of greater interest is the comet of 1138, because two of its components were observed: the 1882 sungrazer (the principal fragment) and Ikeya-Seki. In addition, there are two earlier suspects — X/1702 D1 and the sun-comet of 1792 — and a predicted candidate, expected to arrive within 10 years or so from now. The fragmented nucleus of the 1138 comet is thus modeled on the assumption that it consisted of five condensations. Its length at the time of the comet’s recorded tail sighting, $\sim$33 days after perihelion, is calculated to have amounted to about ${}^{\prime\prime}\!$.
To learn about changes in the overall appearance and morphology that a fragmented nucleus undergoes over a period of time, I present a selected set of four drawings of the nuclear region of the 1882 sungrazer, made by W. L. Elkin at the Cape Observatory between October 13 and November 21, 25 to 64 days after perihelion. All four condensations, for which the positions were calculated from Kreutz’s orbits, are seen only in the first drawing. The separation distances measured by Elkin are in very good agreement with the computed ones.
In the other three drawings fewer than four condensations are apparent, most of them marginally, with the principal fragment B enshrouded in the cloud of debris to the extent that it is completely invisible (with one possible exception). Two drawings are dominated by fragment A, which during October was evidently undergoing a protracted flare-up. By the end of the month it was already fading and nearly gone by November 21.
The 1882 sungrazer was under observation until the end of May 1883, and unlike fragments B and/or C, fragment A was then no longer seen. Its persisting flare-up in October was the kind of event that is often exhibited by short-lived companions of split comets and potentially is a signature of a disintegrating object. It needs to be admitted that there is no correlation between the mass and instantaneous brightness of a sungrazer’s fragment and, as a corollary, that not every bright enough fragment, which is given birth in the process of tidal fragmentation, does necessarily survive throughout the orbit about the Sun to successfully pass — as an essentially intact, sizable sungrazer — yet another perihelion.
This caveat applies to the 1138 comet as much as to the 1882 sungrazer, with one exception: fragments of these two objects cannot readily be compared because of the difference of one generation between them. Because the fragments of the 1882 sungrazer are fragments of a fragment of the 1138 comet, survival throughout the revolution about the Sun is a priori more likely for the fragments of the 1138 comet than the 1882 comet. Still, survival is not warranted for either of them. This uncertainty is a major complication for predictions of future appearances of Kreutz sungrazers that is in principle impossible to bypass.
Concluding this investigation, it is my hope that it has cleared any persisting ambiguity or confusion regarding the parents of the current generation of Populations I and II of the Kreutz system, and once and for all put to rest the lingering problem of a missing second sungrazer.